Grade 12Algebra

cot^-2+cot^-8+cot^-18+_______upto n terms find the sum of this series.

Atul
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
 
firstly in such questions try to find pattern between numbers : here it is 2n and then try to express it in form of tan-1a – tan-1b
∑ cot-1 ( 2n)  = ∑ tan-1 ( 1/2n)   { from n=1 to infinity }
tan-1 ( 2/4n)  =  tan-1 { (2n+1) – (2n - 1) \1 + (2n - 1)(2n+1)  } …......................1
tan-1a - tan-1b = tan-1 {(a-b)/1 + a*b}
so equation 1 equals to tan-1 (2n+1) – tan-1 (2n-1)
 so now the qauestion is ∑ {tan-1 (2n+1) – tan-1 (2n-1) }   [ n from 1 to infinity]
so now the answer is  ( \pi/2 – \pi/4 ) = \pi/4 
 
Regards
Arun